# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 9.1 Subtleties of Reserve Systems
Consider a common scenario in which a decision maker allocates multiple identical, unit-demand, indivisible goods—such as school seats, public jobs, or housing units—based on two criteria. The first criterion is a baseline priority ranking determined by factors like standardized test results, application waiting time, or a lottery. The second criterion adjusts this ranking by giving preferential treatment to specific groups of applicants, such as underrepresented minorities or individuals with particular qualifications. A portion of the available units, the “open” positions, are allocated based solely on the first criterion, while the remaining “reserved” positions are distributed according to the second. This process exemplifies what is called a reserve system.
In this setting, since applicants are priority ranked under both criteria, the implementation might seem straightforward at first. Simply allocate the open positions using the baseline priority ranking and the reserved positions using the modified priority ranking. This seems simple enough. However, as we will show in the next example, there is more to this task than meets the eye.
@@
If all 100 positions were open to everyone, the top 100 scorers would be selected, resulting in exactly 50 men and 50 women being hired—specifically, those who scored 51 or higher. How would the aforementioned positive discrimination policy for women affect this outcome?
Starting with the 70 open positions, let’s allocate all positions based on the specified criteria. Since both groups have the same score distribution, 35 men and 35 women—those with scores of 66 or higher—receive the open positions. The remaining 30 reserved positions are awarded to women with scores between 36 and 65. Overall, men receive 35 positions, and women receive 65, providing women applicants a considerable advantage under this policy.
This solution, however, is not the only way to implement the stated positive discrimination policy. In particular, the outcome would differ significantly if the process began with the allocation of the 30 reserved positions. In that case, the 30 reserved positions would all be awarded to women with scores of 71 and higher. When the 70 open positions are allocated next, women would be at a disadvantage, as the highest-scoring remaining woman has a score of 70. The 30 men with higher scores than any remaining woman would be awarded the first 30 open positions. The playing field is leveled only at that point, with 40 open positions remaining. Since the score distribution of the remaining applicants is identical for men and women at this stage, the last 40 open positions would be evenly split between the two groups, with each receiving 20 positions. Consequently, women would receive only 20 open positions compared to men, who receive 50. Overall, both men and women would receive a total of 50 positions each.
@@
Over-and-Above Choice Rule: The first implementation, which awards open positions before the reserved ones, ensures that the protected group is not disadvantaged in any way during the allocation of open competition positions due to the “cream skimming” of higher-scoring members of the target group for the reserved slots. In a manner of speaking, competition remains truly open in this implementation. First explored by Dur et al. (2018b), the choice rule that implements this form of reserve policy is known as the over-and-above choice rule.
Minimum Guarantee Choice Rule: In contrast, the alternative implementation awards reserved positions before the open ones. This approach significantly disadvantages the beneficiaries of the reserved positions in the allocation of open positions. Aptly called the minimum guarantee choice rule and first explored by Hafalir et al. (2013), this policy merely provides a “minimum guarantee” to the protected group equal to the number of reserved positions. If the group would have already received at least as many positions without positive discrimination—as is the case in Example 7—then they do not benefit at all from this policy. What they gain from the reserved positions, they lose from the open ones.
When reserve systems were first studied by Hafalir et al. (2013) in the market design literature, they presented them directly as the minimum guarantee choice rule,101101101Hafalir et al. (2013) refer to the minimum guarantee choice rule policy as soft reserves. without discussing other possible variants. The significance of the processing sequence of positions—the order of precedence—was first introduced and studied by Kominers and Sönmez (2013, 2016) within a more general framework that incorporates slot-specific priorities into the matching with contracts model discussed in Section 8.4. Building on this use-inspired basic theory, which became pivotal in several subsequent “discovery-invention cycles,” Dur et al. (2018b) examine our current real-world setting as a special case of this model.
As illustrated in Example 7 and formalized in our next result, the over-and-above and minimum guarantee choice rules represent two polar extremes in our setting.
@@
Fix the numbers of open and reserved positions, as well as the order of precedence, and consider the choice rule $C(.)$ that allocates positions sequentially to the highest-priority eligible individual for each position. Then, for any set of applicants:
@@
    The set of protected applicants who receive a position under the minimum guarantee choice rule is a subset of the set of protected applicants who receive a position under the choice rule $C(.)$, which in turn is a subset of the set of protected applicants who receive a position under the over-and-above choice rule.
@@
    The set of unprotected applicants who receive a position under the minimum guarantee choice rule is a subset of the set of unprotected applicants who receive a position under the choice rule $C(.)$, which in turn is a subset of the set of unprotected applicants who receive a position under the over-and-above choice rule.
@@
At first glance, this result may seem counterintuitive, as processing a reserved position earlier might initially appear more advantageous to its beneficiaries. On the contrary, allocating a reserved position earlier exposes the beneficiary group to harsher competition for subsequent open positions, potentially diminishing their overall benefit.
This subtlety in the order of precedence in reserve systems—and particularly the distinction between the over-and-above and minimum guarantee policies—is so elusive that it has led to unintended consequences in several real-world applications. In some cases, decision makers have completely overlooked this distinction, leaving major loopholes in systems that otherwise meticulously control parameters such as the fraction of reserved units.102102102See the The Hindu op-ed Pathak and Sönmez (2019) for an example of such a loophole when the 103rd Constitutional Amendment Act, 2019 introduced reservations for economically weaker sections of society. Occasionally, politically motivated actors have exploited these loopholes through reforms with significant distributional consequences.103103103See Pathak et al. (2025) for an example of this, when officials of the U.S. Department of Homeland Security reversed the processing sequence of two categories of H-1B visas in response to President Trump’s Buy American and Hire American Executive Order in 2017. In other instances, as illustrated in our next case study, system operators intended to implement either the over-and-above or minimum guarantee policy but ended up implementing the wrong one—either because they were unaware of the distinction or did not know how to operationalize the intended policy, especially in a multi-institution setting.
@@ 9.2 The Demise of Neighborhood Priority in Boston School Choice
Over the past fifty years, Boston Public Schools (BPS) have undergone several major reforms in their student assignment policies. From controversial busing mandates to the adoption of state-of-the-art mechanisms developed in collaboration with design economists, the city has consistently aimed to balance equal educational opportunities, racial diversity, and neighborhood priority.
Building on Dur et al. (2018b), our next application examines Boston’s neighborhood priority policy and the unintended consequences of its flawed implementation between 2005 and 2013, which ultimately disadvantaged applicants seeking to attend their neighborhood schools.104104104The period from 2005 to 2013 provides a conservative estimate of the duration of the flawed implementation of Boston’s neighborhood priorities. The neighborhood priority policy intended for use under the DA mechanism during this period was originally adopted in 2000 under Boston mechanism. The flawed implementation of this policy, discussed in Section 9.2.2, led to the de facto elimination of neighborhood priorities under the DA mechanism and would have caused similar unintended consequences in the Boston mechanism had an analogous error occurred. However, it remains unclear whether such a flawed implementation of the neighborhood priority policy existed between 2000 and 2005 under the Boston mechanism.
@@ 9.2.1 Evolution of Neighborhood Priority in Boston
In response to legal challenges to race- and ethnicity-based affirmative action policies across the U.S., the Boston School Committee proactively enacted a major reform in July 1999, eliminating racial and ethnic classifications in Boston’s assignment plan, effective from the 2000-01 school year. Since this reform risked resegregation due to Boston’s demographics, an additional reform was introduced later that year. Crucial to our current discussion, the new system reduced the scope of walk zone priority from 100% to 50% of school seats, balancing neighborhood priority with broader access for all students. Half of the seats at each school were reserved for students within the walk zone, while the other half were open to those outside it.
@@
In media coverage, the 50-50 seat split was often portrayed as a delicate compromise between advocates of neighborhood schools and supporters of broader school choice. Superintendent Payzant expressed hope that the plan would appease both groups by allowing families to send their children to nearby schools while also providing options for those seeking greater choice (Dur et al., 2018b).
Since the city relied on a single uniform lottery to break ties across all schools and both types of seats, achieving this compromise required walk zone priorities to be implemented through the “over-and-above” choice rule. Otherwise, neighborhood students would face a disadvantage when competing for open seats. In the worst-case scenario, as illustrated by Example 7, implementing the intended walk zone priorities via the “minimum guarantee” choice rule would result in the de facto elimination of any positive discrimination for neighborhood students, directly contradicting the spirit of the intended compromise.
Following the publication of Abdulkadiroğlu and Sönmez (2003), BPS adopted the DA mechanism in 2005 after two years of community engagement and an evaluation of the Boston mechanism in collaboration with our team of design economists (see Section 6.10). It remains unclear whether BPS properly implemented its walk zone priority with the “over-and-above” choice rule between 2000 and 2005, prior to this reform. However, it became evident that BPS failed to do so under the DA mechanism starting in 2005, unintentionally leading to the de facto elimination of its positive discrimination policy for neighborhood applicants.
@@ 9.2.2 Unintended Consequences: De Facto Elimination of Boston’s Neighborhood Priority from 2005 to 2013
An unexpected outcome emerged from these counterfactuals: the 50-50 walk zone compromise produced results remarkably similar to those in a system with no walk zone priority at all. How could this happen?
To understand this anomaly, we examined how the city implemented its 50-50 walk zone compromise under the DA mechanism in a multi-institution setting.
Rather than directly embedding a choice rule for each school in the Deferred Acceptance algorithm, the BPS software team implemented the vanilla version of the algorithm by treating each school as two distinct schools with half capacity each and their own priorities—one half with walk zone priority and the other half without. Since students submit preferences over schools rather than their specific halves, the software team artificially constructed preferences for each student $i$ over the “expanded” set of schools:
@@
    For any school $a$, student $i$ prefers the walk zone half to the non-walk zone half.
While this construction may appear natural and without major consequences, the second condition alters the intended neighborhood priority policy. Specifically, this construction meant that students with sufficient priority for both types of seats—for example, those living in the walk zone with favorable lottery numbers—were systematically considered first for the walk zone half. This led to the implementation of the “minimum guarantee” choice rule for each school, rather than the intended “over-and-above” choice rule (Dur et al., 2018b). Consequently, the 50-50 compromise produced results very similar to those in a system with no walk zone priority, significantly shifting the desired balance to the detriment of neighborhood applicants.
@@ 9.2.3 Phasing Out Neighborhood Priority in Boston: 2013 Reform for Transparency
Mayor Thomas Menino’s 2012 address sparked a year-long, city-wide debate on school choice, resulting in numerous proposals from BPS, the EAC, and community members. Over 70 public meetings were held, with input from thousands of parents. A central issue was a puzzling discrepancy: At popular schools, neighborhood students were consistently receiving about 50% of the seats—those reserved for them—but virtually none from the open competition seats. Not only was this highly implausible, but it also undermined the city’s goal of balancing the needs of neighborhood students and providing broader access to all students.
@@
Though I didn’t know it at the time, one of the regular participants in the public meetings was Peng Shi, an MIT graduate student in operations research who was auditing my market design course at Boston College. Recognizing the gridlock as an opportunity, Shi proposed an entirely new approach during one of the meetings: the “Home-Based” system (Shi, 2015).105105105Later, I learned that Shi had been studying the crisis with an eye toward helping guide a reform as early as Spring 2012. Before the Home-Based system, he also proposed other plans when the EAC elicited suggestions from the wider community in late summer. This system provided each student with a customized list of schools based on their home address, reducing the number of schools they could rank while ensuring that highly-rated options were included.106106106See Shi (2022) for an in-depth exploration of Shi’s broader methodology, which optimizes parameters such as student priority distribution within mechanisms like DA and TTC, parameters that are typically treated as exogenous in design economics. His proposal gained traction, as EAC members were eager to bridge the divide between the committee’s opposing camps.
@@
Our discovery of the critical role of the order of precedence—the sequence in which different types of positions are processed—became a central point in the debate between advocates of neighborhood assignment and those supporting increased school choice. Proponents of neighborhood assignment argued that the improper implementation of the 50-50 split had unjustly excluded many students from their neighborhood schools and that correcting the order of precedence was essential to honor the School Committee’s original policy goals from 1999.
@@
In my view, there was also a less benign reason for this faction’s insistence on maintaining the minimum guarantee choice rule: some members of this group appeared tempted to leverage the flawed and misleading implementation of Boston’s neighborhood priority policy—the 50-50 split, viewed as a symbol of compromise—to create the false impression that they were making concessions to neighborhood advocates. While the system emphasized preferential treatment for neighborhood students in half of the seats, it concealed the advantage enjoyed by non-neighborhood students in the other half.

### v4

@@ 9.1 Subtleties of Reserve Systems
Consider a common scenario in which an authority allocates multiple identical indivisible goods with unit demand—such as school seats, public jobs, or housing units—based on two criteria. The first criterion is a baseline priority ranking determined by factors like standardized test results, application waiting time, or a lottery. The second criterion adjusts this ranking by giving preferential treatment to specific groups of applicants, such as underrepresented minorities or individuals with particular qualifications. A portion of the available units—the open positions—are allocated based solely on the first criterion, while the remaining reserved positions are allocated according to the second. This process exemplifies a simple two-category version of what is called a reserve system.
In this setting, since there are two categories of positions with their own priority rankings, the implementation might seem straightforward at first: allocate the open positions using the baseline priority ranking and the reserved positions using the modified priority ranking—a procedure that appears simple enough. However, as we will show in the next example, depicted in Figure 25, there is more to this design exercise than meets the eye.
@@
As a baseline scenario, if all 100 positions were open to everyone, the top 100 scorers would be selected, resulting in exactly 50 men and 50 women being hired—specifically, those who scored 51 or higher. How would the aforementioned positive discrimination policy for women affect this outcome?
Starting with the 70 open positions, suppose we allocate all positions based on the specified criteria. Since both groups have the same score distribution, 35 men and 35 women—those with scores of 66 or higher—receive the open positions. The remaining 30 reserved positions are awarded to women with scores between 36 and 65. Overall, men receive 35 positions, and women receive 65, giving women a considerable advantage under this policy.
This solution, however, is not the only way to implement the stated positive discrimination policy. In particular, the outcome would differ significantly if the process began with the allocation of the 30 reserved positions. In that case, the 30 reserved positions would all be awarded to women with scores of 71 and higher. When the 70 open positions are allocated next, women would be at a disadvantage, as the highest-scoring remaining woman has a score of 70. The 30 men with higher scores than any remaining woman would be awarded the first 30 open positions. The playing field is leveled only at that point, with 40 open positions remaining. Since the score distribution of the remaining applicants is identical for men and women at this stage, the last 40 open positions would be evenly split between the two groups, with each receiving 20 positions. Consequently, women would receive only 20 open positions compared to men, who receive 50. Overall, both men and women would receive 50 positions each.
@@
Figure 25: Merit-based allocation vs. two reserve systems. Panel (a) shows a continuum approximation of the setup: the top blue box represents 100 men (M), and the bottom red box represents 100 women (W), each evenly distributed across test scores from 1 to 100 (with higher scores on the left). The middle empty box represents the 100 available positions (P). Panel (b) illustrates merit-only allocation: all 100 open positions go to the top scorers (scores $\geq 51$), giving 50 positions to men and 50 to women. Panel (c) allocates open positions before reserved ones: the 70 open positions go to the top-scoring individuals, namely 35 men and 35 women (scores $\geq 66$). The 30 reserved positions then go to the next highest-scoring women (scores 65–36), giving 65 positions to women and 35 to men in total. Panel (d) allocates reserved positions before open ones: the first 30 reserved positions go to the highest-scoring women (scores 100–71). With no women left scoring above 70, the next 30 positions, all open, go to men with scores 100–71. The final 40 open positions go to the highest-scoring remaining individuals, split evenly between men and women with scores 70–51, giving 50 positions to women and 50 to men.
@@
Over-and-Above Choice Rule.
The first implementation, which awards open positions before the reserved ones, ensures that the protected group is not disadvantaged in any way during the allocation of open competition positions due to the “cream skimming” of higher-scoring members of the target group for the reserved slots. In a manner of speaking, competition remains truly open in this implementation. First explored by Dur et al. (2018b), the choice rule that implements this form of reserve policy is known as the over-and-above choice rule.
Minimum Guarantee Choice Rule.
In contrast, the alternative implementation awards reserved positions before the open ones. This approach significantly disadvantages the beneficiaries of the reserved positions in the allocation of open positions. Aptly called the minimum guarantee choice rule and first explored by Hafalir et al. (2013), this policy merely provides a “minimum guarantee” to the protected group equal to the number of reserved positions. If the group would have already received at least as many positions without positive discrimination—as is the case in Example 11—then they do not benefit at all from this policy. What they gain from the reserved positions, they lose from the open ones.
In Figure 25, Panels (c) and (d) depict the workings of the over-and-above and minimum guarantee choice rules, respectively.
When reserve systems were first studied by Hafalir et al. (2013) in the market design literature, they were presented directly as the minimum guarantee choice rule,113113113Hafalir et al. (2013) refer to the minimum guarantee choice rule policy as soft reserves. without discussion of other possible variants. The significance of the processing sequence of positions—the order of precedence—was first introduced and studied by Kominers and Sönmez (2013, 2016) within a more general framework that incorporates slot-specific priorities into the matching-with-contracts model discussed in Section 8.4. Building on this use-inspired basic theory—pivotal in several subsequent “discovery—invention cycles”—Dur et al. (2018b) examine the simpler setting discussed in this section, which is widespread in real-world applications, as a special case of that model (see Figure 36).
As illustrated in Example 11 and formalized in our next result, the over-and-above and minimum guarantee choice rules represent two polar extremes in our current setting.
@@
Suppose there are two types of positions—open and reserved—that are allocated based on a baseline priority order subject to eligibility. Let there be a set of applicants who are all eligible for the open positions, and a single protected subset of these applicants who are exclusively eligible for the reserved positions. Fix the numbers of open and reserved positions, as well as the order of precedence, and consider the choice rule $C(\cdot)$, which allocates positions sequentially to the highest-priority eligible individual for each position. Then:
@@
    The set of protected applicants who receive a position under the minimum-guarantee choice rule is a subset of the set of protected applicants who receive a position under the choice rule $C(\cdot)$, which in turn is a subset of the set of protected applicants who receive a position under the over-and-above choice rule.
@@
    The set of unprotected applicants who receive a position under the over-and-above choice rule is a subset of the set of unprotected applicants who receive a position under the choice rule $C(\cdot)$, which in turn is a subset of the set of unprotected applicants who receive a position under the minimum-guarantee choice rule.
@@
At first glance, this result may seem counterintuitive, since processing a reserved position earlier might appear more advantageous to its beneficiaries. Yet, in reality, allocating a reserved position earlier exposes the beneficiary group to harsher competition for subsequent open positions, potentially reducing their overall benefit. Panel (d) of Figure 25 illustrates the worst-case scenario—under the minimum guarantee choice rule, where all reserved positions are processed before all open ones—in which the beneficiary group loses all advantage relative to pure merit-based allocation.114114114In general, however, the loss may be less severe, with the group retaining part of the benefit when pure merit-based allocation falls short of the minimum guarantee.
This subtlety in the order of precedence in reserve systems—particularly the distinction between the over-and-above and minimum guarantee policies—is so elusive that it has led to unintended consequences in several real-world applications. In some cases, decision-makers have completely overlooked this distinction, leaving major loopholes in systems that otherwise meticulously stipulated more straightforward parameters, such as the fraction of reserved units.115115115See the The Hindu op-ed Pathak and Sönmez (2019) for an example of such a loophole, when the 103rd Constitutional Amendment Act, 2019 introduced reservations for economically weaker sections of society. Occasionally, politically motivated parties have leveraged these loopholes through reforms with significant distributional consequences.116116116See Pathak et al. (2025) for an example of this, when officials of the U.S. Department of Homeland Security reversed the processing sequence of two categories of H-1B visas in response to President Trump’s Buy American and Hire American Executive Order in 2017. In other instances, as illustrated in our next case study, system operators intended to implement either the over-and-above or minimum guarantee policy but ended up implementing the wrong one—either because they were unaware of the distinction or did not know how to operationalize the intended policy, especially in a multi-institution setting.
@@ 9.2 The Demise of Neighborhood Priority in Boston School Choice
Over the past fifty years, Boston Public Schools (BPS) have undergone several major reforms in their student assignment policies. From controversial busing mandates to the adoption of state-of-the-art mechanisms developed in collaboration with economic designers, the city has consistently aimed to balance equal educational opportunities, racial diversity, and neighborhood priority.
Building on Dur et al. (2018b), our next application examines Boston’s neighborhood priority policy and the unintended consequences of its flawed implementation between 2005 and 2013, which ultimately disadvantaged applicants seeking to attend their neighborhood schools.117117117The period from 2005 to 2013 provides a conservative estimate of the duration of the flawed implementation of Boston’s neighborhood priorities. The neighborhood priority policy, intended for use under the DA mechanism during this period, was originally adopted in 2000 under the Boston mechanism. The flawed implementation of this policy, discussed in Section 9.2.2, led to the de facto elimination of neighborhood priorities under the DA mechanism and would have caused similar unintended consequences under the Boston mechanism had an analogous error occurred, though whether such a flaw existed between 2000 and 2005 remains unclear.
@@ 9.2.1 Evolution of Neighborhood Priority in Boston
In response to legal challenges to race- and ethnicity-based affirmative action policies across the U.S., the Boston School Committee proactively enacted a major reform in July 1999, eliminating racial and ethnic classifications in Boston’s assignment plan, effective from the 2000–01 school year. Since this reform risked resegregation due to Boston’s demographics, an additional reform was introduced later that year. Crucial to our current discussion, the new system reduced the scope of walk zone priority from 100% to 50% of school seats, balancing neighborhood priority with broader access for all students. Half of the seats at each school were reserved for students within the walk zone, while the other half were open to students from outside the zone.
@@
In media coverage, the 50–50 seat split was often portrayed as a delicate compromise between advocates of neighborhood assignment and supporters of broader school choice. Superintendent Payzant expressed hope that the plan would appease both groups by allowing families to send their children to nearby schools while also providing options for families seeking greater choice (Dur et al., 2018b).
Since the city relied on a single uniform lottery to break ties across all schools and both types of seats, achieving this compromise required implementing walk zone priorities through the over-and-above choice rule (Panel (c) in Figure 25). Otherwise, neighborhood students would face a disadvantage when competing for open seats. In the worst-case scenario, as shown in Example 11, implementing the intended walk zone priorities via the minimum guarantee choice rule would effectively eliminate any positive discrimination for neighborhood students, directly contradicting the spirit of the intended compromise (Panel (d) in Figure 25).
Following the publication of Abdulkadiroğlu and Sönmez (2003), BPS adopted the DA mechanism in 2005 after two years of community engagement and an evaluation of the Boston mechanism in collaboration with our team of economic designers (see Section 6.10). It remains unclear whether BPS properly implemented its walk zone priority with the over-and-above choice rule between 2000 and 2005, prior to this reform. However, it became evident that BPS failed to do so under the DA mechanism starting in 2005, unintentionally leading to the de facto elimination of its positive discrimination policy for neighborhood applicants.
@@ 9.2.2 Unintended Consequences: De Facto Elimination of Boston’s Neighborhood Priority from 2005 to 2013
An unexpected outcome emerged from these counterfactuals: the 50–50 walk zone compromise produced results remarkably similar to those in a system with no walk zone priority at all. How could this happen?
To understand this anomaly, we examined how the city implemented its 50–50 walk-zone compromise under the DA mechanism in a multi-institution setting.
Rather than embedding a choice rule that reflected the city’s intended compromise policy directly into the deferred acceptance algorithm, the BPS software team implemented the vanilla version by treating each school as two distinct schools, each with half capacity and its own priorities—one with walk-zone priority and the other without. Even though sibling priority also applied for both halves, to simplify the discussion, let’s refer to them as the walk-zone half and the open-competition half.
Since students submit preferences over schools rather than their specific halves, the software team artificially constructed preferences for each student $i$ over the “expanded” set of schools:
@@
    For any school $a$, student $i$ prefers the walk-zone half to the open-competition half.
While this construction may have appeared natural and without major consequences, the second condition in fact altered the intended neighborhood priority policy. Specifically, it meant that students with sufficient priority for both types of seats—for example, those living in the walk zone with favorable lottery numbers—were systematically considered first for the walk-zone half. This led to the implementation of the minimum guarantee choice rule for each school, rather than the intended over-and-above choice rule (Dur et al., 2018b). Intuitively, when the open-competition half was considered at a given school, walk-zone applicants systematically had less favorable lottery numbers, since those with better numbers had already been considered for seats in the walk-zone half. Consequently, the 50–50 compromise produced outcomes very similar to those in a system with no walk-zone priority, significantly shifting the balance to the detriment of neighborhood applicants.
@@ 9.2.3 Phasing Out Neighborhood Priority in Boston: 2013 Reform for Transparency
Mayor Thomas Menino’s 2012 address sparked a year-long, city-wide debate on school choice, resulting in numerous proposals from BPS, the EAC, and community members. Over 70 public meetings were held, with input from thousands of parents. A central issue was a puzzling discrepancy: At popular schools, neighborhood students were consistently receiving about 50% of the seats—those reserved for them—but virtually none from the open-competition seats. Not only was this highly implausible, but it also undermined the city’s goal of balancing the needs of neighborhood students and providing broader access to all students.
@@
Though I didn’t know it at the time, one of the regular participants in the public meetings was Peng Shi, an MIT graduate student in operations research who was auditing my market design course at Boston College. Recognizing the gridlock as an opportunity, Shi proposed an entirely new approach during one of the meetings: the “Home-Based” system (Shi, 2015).118118118I later learned that Shi had been studying the crisis with the aim of informing a reform as early as the spring of 2012. Before the adoption of the Home-Based system, he developed several alternative proposals, including one—later published in Ashlagi and Shi (2014)—for improving community cohesion in school choice through a correlated-lottery implementation that generalized earlier work by Liu and Pycia (2016). Some of these proposals were submitted when the EAC invited suggestions from the wider community in late summer. This system provided each student with a customized list of schools based on their home address, reducing the number of schools they could rank while ensuring that highly-rated options were included.119119119See Shi (2022) for an in-depth exploration of Shi’s broader methodology, which optimizes parameters such as student priority distribution within mechanisms like DA and SC–TTC, parameters that are typically treated as exogenous in economic design. His proposal gained traction, as EAC members were eager to bridge the divide between the committee’s opposing camps.
@@
Our discovery of the critical role of the order of precedence—the sequence in which different types of positions are processed—became a central point in the debate between advocates of neighborhood assignment and those supporting increased school choice. Proponents of neighborhood assignment argued that the improper implementation of the 50–50 split had unjustly excluded many students from their neighborhood schools and that correcting the order of precedence was essential to honor the School Committee’s original policy goals from 1999.
@@
In my view, there was also a less benign reason for this faction’s insistence on maintaining the minimum guarantee choice rule: some members of this group appeared tempted to leverage the flawed and misleading implementation of Boston’s neighborhood priority policy—the 50–50 split, viewed as a symbol of compromise—to create the false impression that they were making concessions to neighborhood advocates. While the system emphasized preferential treatment for neighborhood students in half of the seats, it concealed the advantage enjoyed by non-neighborhood students in the other half.